Find AI, IBI, AB, and ABI. Then verify that A||B| = |ABI. [012] [ 2 1 2] 5 4 3. B 1 -1 0 | 7 6 8 09-2 (a) ALL (b) 18 (d) ABPlease help with i
Answer #1

rol 2) Is 4 3 17 68 lala o ri(40-21 ) 42 (80-28) - 19 + 4 = miš 16 10) = -15 1 B = {2 127 lo 9-2 131 = CQ-0)-1(~29+26) = 4+2+@ 1AB) = 1.(-56 - 292)-17 (-28-80).-4 (1022-560) = -348 +1836 – 1848 – – 860 I IABI=-360